Please answer both questions 3 and 4. Thank you!3) The half-life for the 1st order decay of a neurotoxin is found to be 2.5 hrs. If a patient reports to have ingested a init
Answer #1

Question number (3)

We know that for first Order reaction.

k = 0.693/t​​​​​​1/2 .........(1)

t = (2.303/k)log(a/a-x). .......(2)

Where,

k = rate constant

t​​​​​​1/2 = half life period = 2.5hrs

t = total time

a = initial consultation of reactant = 200microMolar

a-x = final concentration of reactant = 1micromolar

Putting the value of k in equation number (2)

t = (2.303t1/2/0.693)log(a/a-x)

Using the given values

t = (2.303×2.5hrs/0.693)log(200/1)

t = (8.31hrs)×log(200)

t = (8.31hrs)(2.3010)

t = 19.1hrs

Question (4) we know that,

number of mole of X = (concentration of x)(volume of X)

1000mL = 1L

1M = 1mol/L

Step(1) Calculation For Number of mole of sulfuric acid.

Volume of sulfuric acid = 100mL = 0.1L

Concentration of sulfuric acid = 0.100M = 0.100mol/L

Using the formula

Number of mole of Sulfuric acid = (0.100mol/L)(0.1L)

Number of mole of Sulfuric acid = 0.010mol

Step(2) Calculation For number of mole of Barium hydroxide.

Volume of Ba(OH)2 = 200mL = 0.2L

Concentration of Ba(OH)2 = 0.15M = 0.15mol/L

Using the formula

Number of mole of Ba(OH)2 = (0.15mol/L)(0.2L)

Number of mole of Ba(OH)2 = 0.03mol

Step(3) Calculation For Unreacted number of mole of Ba(OH)2 .

Chemical reaction between sulfuric acid and Barium hydroxide will be given as fofollows.

H2SO​​​​4  + Ba(OH)2 --------> Ba(SO)4  + 2H2O

From the stoichiometry of chemical reaction.

1mol sulfuric acid completely reacted with 1mol Barium hydroxide.

So

0.01mol sulfuric acid completely reacted with 0.01mol Barium hydroxide.

Unreacted Barium hydroxide = ( total Barium hydroxide) - (reacted Barium hydroxide)

Unreacted Barium hydroxide = 0.03mol - 0.01mol

Unreacted Barium hydroxide = 0.02mol

Step(4) Calculation For number of mole of OH​​​​​-

Ba(OH)2  ---------------> Ba​​​​​2+  + 2OH-

From the stoichiometry of chemical reaction.

1mole Barium hydroxide gives 2 mole of OH​​​​​- .

So

0.02mole Barium hydroxide given 0.04mol of OH​​​-

Number of mole of OH​​​​​- = 0.04mol

Step(5) Calculation For concentration of OH​​​​​-

Concentration = (number of mole )/(total volume)

Total volume = volume of sulfuric acid + volume of Barium hydroxide

Total volume = 100mL + 200mL = 300mL = 0.3L

So,

Concentration of OH​​​​​- = [OH​​​​​- ] = (0.04mol)/(0.3L)

[OH​​​​​-​] = 0.012mol/L = 0.012M

Step(6) Calculation For P​​​​​​H .

We know that,

P​​​​​​H = -log[H+ ] .......(1)

Or

P​​​​​​H = 14 + log[OH​​​​​- ] .......(2)

Using the equation number (2)

P​​​​​​H = 14 + log(0.012)

P​​​​​​H = 14 - 1.92

P​​​​​​H = 12.08